# Deep Generative Classification of Blood Cell Morphology

### v1

@@ Abstract
Accurate classification of haematological cells is critical for diagnosing blood disorders, but presents significant challenges for machine automation owing to the complexity of cell morphology, heterogeneities of biological, pathological, and imaging characteristics, and the imbalance of cell type frequencies. We introduce CytoDiffusion, a diffusion-based classifier that effectively models blood cell morphology, combining accurate classification with robust anomaly detection, resistance to distributional shifts, interpretability, data efficiency, and superhuman uncertainty quantification. Our approach outperforms state-of-the-art discriminative models in anomaly detection (AUC 0.976 vs. 0.919), resistance to domain shifts (85.85% vs. 74.38% balanced accuracy), and performance in low-data regimes (95.88% vs. 94.95% balanced accuracy). Notably, our model generates synthetic blood cell images that are nearly indistinguishable from real images, as demonstrated by a Turing test in which expert haematologists achieved only 52.3% accuracy (95% CI: [50.5%, 54.2%]). Furthermore, we enhance model explainability through the generation of directly interpretable counterfactual heatmaps. Our comprehensive evaluation framework, encompassing these multiple performance dimensions, establishes a new benchmark for medical image analysis in haematology, ultimately enabling improved diagnostic accuracy in clinical settings. Our code is available at https://github.com/Deltadahl/CytoDiffusion.
@@
8Queen Mary University of London, London, United Kingdom
@@
The haematological is amongst the most complex of physiological systems, and uniquely intertwined with all others. Though often quantified by simple “blood counts” of cell class frequencies, its characteristics are both supremely rich and highly variable within and across individuals [1]. A cardinal aspect is the morphology of individual blood cells as seen on light microscopy of blood smears [2]. The richness of morphological appearances—and their complex modulation by diverse biological, pathological, and instrumental factors—demand human expert visual description in commensurately expressive terms. The allocation of a cell to its major morphological type, e.g. lymphocyte, is only the crudest form of description, on which finer fractionation into subtypes, across a wide spectrum of (ab)normality, is overlaid.
Indeed, the task of morphological characterisation is both open-ended and lacks a definitive ground truth: there may be morphological patterns whose subtlety has concealed great clinical significance, and some morphological classes are purely expert-determined visual phenotypes with no means of objective corroboration. Moreover, pathological appearances may be highly unusual or unique, precluding classification into any class, even at the simplest level of description, and requiring metacognitive awareness of its impossibility. The difficulty is commonly compounded by interactions with irrelevant biological features with variable representation across the population, and instrumental variations of technical origin [3, 4]. The challenge, in short, is one human experts can only imperfectly meet, inevitably exhibiting marked variation with skill and experience [5, 6].
@@
Here we therefore introduce a deep generative approach founded on diffusion-based classifiers, and a comprehensive evaluative framework that aspires to real-world ecological validity. Our approach, which we call CytoDiffusion, addresses the key challenges, exemplified by blood morphological analysis but not unique to it, of machine vision models of medical data, namely:
  1. 1.
    Domain Shift: Models trained on data from a specific biological, pathological, and instrumental context often struggle when applied to data with a different distribution of characteristics, limiting their practical utility [7, 8].
  2. 2.
    Anomaly Detection: The ability to identify rare or previously unseen patterns of features is crucial in clinical applications yet often overlooked in model development and evaluation [9]. Models should be capable of recognising patterns absent from the training data, for such cases fall outside the model’s competence and must be so highlighted.
  3. 3.
    Uncertainty Quantification: Ideal epistemic agents–human or machine–can accurately quantify the reliability of their knowledge, exhibiting metacognitive awareness. Reliable uncertainty quantification is crucial for clinical decision-making, yet is widely neglected in model assessments [10]. The utility of a prediction depends on its uncertainty, and requires weighting by it, both in direct use and the design of ensemble-based systems.
  4. 4.
    Data Heterogeneity and Incompleteness: Many existing medical datasets fail to replicate the heterogeneity—biological, pathological, and instrumental—and incompleteness prevalent in real-world clinical life [11]. Generalisation to real-world scenarios is therefore inadequately tested.
  5. 5.
    Interpretability: In medical applications, the reasoning behind a model’s decisions may be as important as the decisions themselves, yet many current approaches lack interpretability [12, 13].
  6. 6.
    Data efficiency: The medical domain is constitutionally constrained in data scale, in numbers of cases in proportion to observed diversity and especially in ground truth labels. Successfully deployable models must therefore be data efficient.
While numerous studies have reported high accuracies in various medical image analysis tasks [14, 15, 16], these broader challenges remain largely unaddressed, limiting the clinical applicability of such models.
We apply diffusion-based classifiers to this real-world clinical challenge, building on recent work that has demonstrated their promising robustness to distribution shifts and ability to avoid shortcut solutions [17, 18, 19, 20]. CytoDiffusion offers several advantages, including improved robustness to domain shifts and inherent capabilities for anomaly detection and uncertainty quantification.
A key advantage of generative classifiers is their ability to model the input in its entirety, rather than relying on spurious correlations or shortcuts that traditional discriminative classifiers often exploit [17]. By learning to generate examples of each class, generative classifiers are forced to capture the full complexity of the input space.
In the context of blood cell classification, CytoDiffusion is compelled to learn the complete morphological characteristics of each cell type, rather than focusing on easily discriminative but potentially unreliable features. This holistic approach is particularly valuable in medical imaging, where subtle details can be crucial for accurate diagnosis. Figure 1 illustrates the operation of our proposed model.
Specifically, our contributions include:
  1. 1.
    A novel application of latent diffusion models for blood cell classification.
  2. 2.
    An evaluation framework that goes beyond accuracy, incorporating domain shift robustness, anomaly detection capability, and performance in low-data regimes.
  3. 3.
    A new dataset of blood cell images that includes artefacts and labeller confidence scores, addressing key limitations in existing datasets.
  4. 4.
    A principled framework for metacognitive evaluation of model and human confidence based on established psychometric modelling techniques.
  5. 5.
    A method for generating interpretable heatmaps to explain model decisions, crucial for validation.
Through these contributions, we aim to establish a new standard for the development and assessment of blood cell classification models. Our work addresses critical aspects of clinical applicability, including robustness, interpretability, and reliability.
@@
We propose that the research community adopt these evaluation tasks and metrics when assessing new models for blood cell image classification. By going beyond simple accuracy measures, we can develop models that are not only high-performing but also trustworthy and clinically relevant.
1 Results
@@
We begin by validating the quality of images generated by CytoDiffusion through a Turing test. Next, we assess the model’s in-domain performance on standard classification tasks across multiple datasets. We then examine CytoDiffusion’s ability to quantify uncertainty, comparing its metacognitive capabilities with those of human experts. Following this, we evaluate the model’s proficiency in anomaly detection, crucial for identifying rare or unseen cell types. We proceed to test the model’s robustness to domain shifts, simulating real-world variability in imaging conditions. Subsequently, we investigate CytoDiffusion’s efficiency in low-data regimes, a critical consideration for medical applications where large, well-annotated datasets may be scarce. Finally, we demonstrate CytoDiffusion’s explainability through the generation of counterfactual heatmaps, providing interpretable insights into its decision-making process.
1.1 Turing test
To assess whether our fine-tuned diffusion model could effectively capture the distribution of blood cell images and demonstrate sufficient expressiveness, we designed a Turing test specifically adapted for peripheral blood film imaging. CytoDiffusion was trained on a dataset comprising 32,619 images. The results of our Turing test, which involved expert haematologists assessing the authenticity of the images, underscore the model’s capability to replicate the complex morphological characteristics of blood cells.
@@ 1.1 Turing test
The 10 haematologists participating in the study evaluated a total of 288 images each, resulting in 2,880 individual assessments. They demonstrated an overall accuracy of 52.3%, 95% CI: $[50.5\%,54.2\%]$ in distinguishing between real and synthetic images, with a sensitivity of 55.8% and a specificity of 48.9%. This performance is comparable to random guessing, indicating that the synthetic images produced by CytoDiffusion are almost indistinguishable from real blood cell images, even to experienced professionals.
@@
In addition to the Turing test, we evaluated the fidelity of blood cell type conditional synthesis. The haematologists’ classification of the synthetic images aligned with the model’s intended cell type with an agreement rate of 98.6%. Overall, these results affirm that CytoDiffusion not only generates high-quality synthetic images but also accurately captures the class-defining morphological features of blood cells (see Figure 8 in the supplementary material).
1.2 In-Domain Performance
@@
To establish a baseline for our diffusion-based classifier, we evaluated CytoDiffusion’s performance on standard in-domain classification tasks. We used three datasets for this evaluation: our custom dataset, Raabin-WBC [21], and PBC [22]. Table 1 presents a comparison of our model against other methods that have been evaluated on these datasets.
CytoDiffusion demonstrates competitive performance across all datasets, particularly in balanced accuracy. This metric is especially relevant given the significant class imbalances present in these datasets. For instance, our custom dataset’s class distribution ranges from 103 to 1425 images per class, whilst Raabin-WBC TestA varies from 89 to 2660 images per class.
It is important to note that we have excluded some studies from our comparison due to methodological differences that could lead to unfair or misleading comparisons. Specifically, we are not comparing with papers that do not have a proper train–validation–test split [23, 24], or those that do not test on the predefined test set [16, 25].
Table 1: Comparison of various methods on our custom dataset, the Raabin-WBC (TestA) dataset, and the PBC dataset. Best results are in bold, second-best are underlined. Balanced accuracy is reported where available.
@@
Raabin-WBC TestA Tavakoli et al. [26] 2021 94.65 92.76
@@
1.3 Uncertainty Measure
@@
Our dataset uniquely incorporates labeller confidence for all images, providing a rare opportunity to compare model uncertainty with human expert uncertainty. This feature allows us to assess how well CytoDiffusion’s uncertainty aligns with that of experienced haematologists, offering valuable insights into the model’s ability to capture the nuances and complexities of cell classification. CytoDiffusion determines the error between the predicted and true noise for each class (see Figure 1). We take the difference between the two smallest errors to represent the confidence of the model: the larger the difference, the more confident the model is in its prediction.
An ideal agent’s epistemic uncertainty is equal to the system’s aleatoric uncertainty. The optimal relation between uncertainty and accuracy should therefore approximate that of an ideal psychophysical observer detecting a noisy signal [40]. Though we do not have independent quantification of the signal here, we can quantify how well the uncertainty measure conforms to a psychophysical model by treating it as if it were an index of task difficulty. We do this by employing Bayesian psychometric function estimation [41, 42] with the inverse of uncertainty–confidence–as the index of task difficulty, estimating the parameters of the resultant psychometric function for both individual experts and our model (Figure 2). Strikingly, whilst expert performance produces a noisy function (Figure 2C), CytoDiffusion conforms to the behaviour of a psychophysical observer with tight bounds on the posterior distributions of the key parameters of width and threshold (Figure 2A). This behaviour closely resembles that of an ideal observer, suggesting that our model’s epistemic uncertainty conforms to the structure of the aleatoric floor.
Indeed, CytoDiffusion produces arguably superior stratification of case difficulty than human experts themselves. Estimating functions of expert performance with model confidence as the index yields tighter fits, and enables better separation of experts differing in performance relative to the consensus ground truth (Figure 2B, D, E).
@@
Figure 2: Bayesian psychometric analysis of model and expert performance. A. Psychometric function of CytoDiffusion’s performance on our custom dataset with model confidence as the index of signal strength. The 95% credibility interval of the threshold of the function, estimated at the point of 80% correct unscaled by lapse and guess rate, is shown as an error bar, and the posterior distributions of threshold and width are presented in the inset axes. Note tight bounds on the parameters. B. Psychometric function of the performance of a single human expert, derived from a randomly selected subset of 200 cases for which a consensus ground truth is available, with model confidence as the index of signal strength. C. Psychometric function of the expert shown in B, with consensus expert confidence as the index of signal strength. D & E. Width and threshold parameter estimates and their 95% credibility intervals for each of the six human experts with model confidence (D) and consensus expert confidence (E) as the indices of signal strength.
CytoDiffusion’s ability to quantify its own uncertainty accurately is crucial for clinical applications, as it allows for the identification of cases that may require additional expert review. This feature could significantly enhance the reliability and trustworthiness of automated blood cell classification systems in real-world clinical settings.
1.4 Anomaly Detection
To assess the model’s ability to distinguish between seen and unseen cell types, we require a metacognitive index of its confidence. The evaluation procedure involved two steps: first, training CytoDiffusion on datasets that omitted specific cell types; then, testing it on held-out sets that included these previously unseen cell types.
To quantify the model’s confidence, we use the measure described in Section 1.3. For normal cells (those present in the training data), we expect the model to be more confident in its top prediction, whilst for abnormal cells (those excluded from training), we expect less confidence.
We visualise these differences for normal and abnormal cells using kernel density estimates (KDE) in Figure 3, revealing a striking difference between the two distributions. To quantify this separation, we constructed receiver operating characteristic (ROC) curves for distinguishing normal from abnormal cells using the error difference. We then calculated the area under the ROC curve (AUC) to measure the model’s ability to differentiate between known and unknown cell types.
@@ 1.4 Anomaly Detection
For the Bodzas dataset [43], with blasts as the abnormal class, we achieved an AUC of 0.971. The PBC dataset [22], using erythroblasts as abnormal, yielded an even higher AUC of 0.976. These high AUC values demonstrate our model’s capability to distinguish between normal cells it was trained on and abnormal cell types not present in the training data.
To put our results in perspective, we compared CytoDiffusion’s performance with a Vision Transformer (ViT-B/16) model. For the ViT, we used the difference between the two largest outputs in the logits (before the softmax) as a measure of confidence. The ViT achieved an AUC of 0.744 for the Bodzas dataset (blasts as abnormal) and 0.919 for the PBC dataset (erythroblasts as abnormal). While the ViT demonstrates reasonable performance, particularly for the PBC dataset, our diffusion-based model consistently outperforms it, showing superior capability in detecting anomalous cell types.
This ability to identify rare or unexpected cell types is crucial in clinical settings, where the detection of abnormal cells can have significant diagnostic implications. We propose that future models report AUC scores for similar anomaly detection tasks, providing a standardised metric for assessing a model’s ability to flag potentially abnormal findings in medical imaging.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionIncreasing certaintyKernel Density EstimateCytoDiffusion (ours)ViT-B/16BlastErythroblast
@@
Figure 3: Kernel density estimate figures comparing the anomaly detection performance of our diffusion-based model (top row) with ViT-B/16 (bottom row) for blasts (left column) and erythroblasts (right column). The horizontal axis represents the certainty measure, normalised to $[0,1]$ by dividing by the maximum certainty for each model and dataset. The AUC values indicate the model’s ability to distinguish between normal and abnormal samples.
1.5 Domain Shift
The ability of a classification model to generalise across different imaging conditions is crucial for its practical application in clinical settings. This generalisation capability is particularly important in haematology, where variations in microscope types, camera systems, and staining techniques are common across different laboratories and hospitals. A model that performs well only under specific imaging conditions would have limited utility in real-world clinical practice, potentially leading to inconsistent or unreliable diagnoses when deployed in new environments.
We propose that new models in this domain should be evaluated on their performance under various domain shift scenarios to assess their robustness and adaptability. To demonstrate this evaluation approach, we conducted experiments using the Raabin-WBC dataset for training, with testing performed on two distinct datasets: TestB and LISC (see Table 3 for dataset details).
TestB is part of the Raabin dataset, but the authors have provided a specific split for testing purposes. It differs from the training set in camera type and microscope type. LISC, on the other hand, introduces more significant shifts, differing in camera type, microscope type, and staining technique. Table 2 presents the results of our evaluation using this framework.
Table 2: Performance under domain shift conditions. Models were trained on the Raabin-WBC dataset and evaluated on TestB and LISC datasets, representing different degrees of domain shift. TestB differs from the training set in camera and microscope type, while LISC introduces additional variations in staining technique. Balanced accuracy scores (mean $\pm$ standard deviation) from five independent training sessions are reported.
Method TestB LISC
@@ 1.5 Domain Shift
1.6 Efficiency in Low-Data Regimes
The ability of a model to perform well with limited training data is crucial in medical imaging contexts, where large, well-annotated datasets are often challenging to obtain. We propose that new models in this domain should be evaluated on their performance in low-data regimes to assess their efficiency.
To demonstrate this evaluation approach, we conducted experiments using the PBC dataset [22]. We compared the performance of multiple models, including our proposed approach CytoDiffusion, and two established architectures: EfficientNetV2-M [45] and ViT-B/16 [44]. The training was carried out with subsets of 10, 20, 50, and 150 images per class, simulating conditions of sparse data availability.
Figure 4 presents the results of these experiments, illustrating the balanced accuracy of each model as a function of the number of training images per class. Our proposed model consistently outperforms EfficientNetV2-M and ViT-B/16 across all data regimes, with the advantage particularly pronounced in the most data-scarce conditions.
@@ 1.6 Efficiency in Low-Data Regimes
Figure 4: Model performance comparison under low-data conditions. The graph shows the balanced accuracy of our proposed model, EfficientNetV2-M, and ViT-B/16 as a function of the number of training images per class. Shaded areas represent the standard deviation from five independent training sessions.
1.7 Explainability
A crucial aspect of deploying machine learning models in clinical settings is their ability to provide interpretable results. CytoDiffusion addresses this need through the generation of counterfactual heatmaps, which highlight the regions of an image that would need to change for it to be classified as a different cell type. A distinctive feature of our approach is that these heatmaps are the direct output of our model, requiring no post-processing or additional colour mapping. This direct output capability preserves the integrity of the model’s decision-making rationale, offering an unfiltered view of its reasoning process.
@@ 1.7 Explainability
Figure 5 illustrates this process using an eosinophil as an example. When the model is prompted to consider what alterations would be necessary for this cell to be classified as a neutrophil, it generates a heatmap ($\boldsymbol{H}_{\text{neutrophil}}$) highlighting the key areas of difference. The overlay of this heatmap on the original image reveals that the model focuses primarily on distinguishing granularity between neutrophils and eosinophils. Eosinophils contain larger, spherical and mostly higher quantity of eosinophilic stained granules in comparison to neutrophils [46].
@@
Refer to captionOriginal “If this image were a neutrophil, what would you change?” Refer to captionHeatmapRefer to captionOverlay
@@
Figure 5: Counterfactual heatmap visualisation. Left: Original eosinophil image. Centre-right: Counterfactual heatmap ($\boldsymbol{H}_{\text{neutrophil}}$) illustrating the areas that differ when the model considers the image as a neutrophil. Darker areas in the heatmaps indicate features to be added, whilst lighter areas suggest features to be removed to achieve the target cell type. Far right: Overlay of the thresholded heatmap on the original image, highlighting the most significant areas of difference, primarily in the granules.
@@
To provide a comprehensive view of our model’s reasoning across all cell types, we generated counterfactual heatmaps for each possible class transition in the PBC dataset (Figure 6). This visualisation offers insights into the model’s decision-making process for each cell type. For instance, when considering the transition from neutrophil to eosinophil (row 2, column 7), the model highlights regions in the cytoplasm (darker areas) where features should be added, while largely maintaining the nuclear shape.
@@
These results provide valuable insights into CytoDiffusion’s decision-making process, offering a detailed visualisation of the morphological distinctions the model identifies between cell types. This granular view of the model’s reasoning enhances our understanding of its classificatory mechanisms, potentially contributing to the development of more transparent and interpretable models across various domains of medical image analysis.

### v2

@@ Abstract
Accurate classification of haematological cells is critical for diagnosing blood disorders, but presents significant challenges for machine automation owing to the complexity of cell morphology, heterogeneities of biological, pathological, and imaging characteristics, and the imbalance of cell type frequencies. We introduce CytoDiffusion, a diffusion-based classifier that effectively models blood cell morphology, combining accurate classification with robust anomaly detection, resistance to distributional shifts, interpretability, data efficiency, and superhuman uncertainty quantification. Our approach outperforms state-of-the-art discriminative models in anomaly detection (AUC 0.990 vs. 0.918), resistance to domain shifts (85.85% vs. 74.38% balanced accuracy), and performance in low-data regimes (95.88% vs. 94.95% balanced accuracy). Notably, our model generates synthetic blood cell images that are nearly indistinguishable from real images, as demonstrated by an authenticity test in which expert haematologists achieved only 52.3% accuracy (95% CI: [50.5%, 54.2%]) in distinguishing real from generated images. Furthermore, we enhance model explainability through the generation of directly interpretable counterfactual heatmaps. Our comprehensive evaluation framework, encompassing these multiple performance dimensions, establishes a new benchmark for medical image analysis in haematology, ultimately enabling improved diagnostic accuracy in clinical settings. Our code is available at https://github.com/CambridgeCIA/CytoDiffusion.
@@
8Queen Mary University of London, London, UK
@@
The haematological is amongst the most complex of physiological systems, and uniquely intertwined with all others. Though often quantified by simple “blood counts” of cell class frequencies, its characteristics are both supremely rich and highly variable within and across individuals [1]. Characterising the morphological appearances of individual blood cells as seen on light microscopy is often critical to managing haematological disorders. Complex modulation of cell morphology by diverse biological, pathological, and instrumental factors, requires that this task necessarily be performed by trained experts. Moreover, labelling cells by their major morphological type, e.g. lymphocyte, is only the crudest form of description, on which finer fractionation into subtypes, across a wide spectrum of (ab)normality, is overlaid.
Indeed, the task of morphological characterisation is both open-ended and lacks a definitive ground truth: there may be morphological patterns whose subtlety has concealed great clinical significance, and some morphological classes are purely expert-determined visual phenotypes with no means of objective corroboration. Moreover, pathological appearances may be highly unusual or unique, precluding classification into any class, even at the simplest level of description, and such anomalies should be explicitly identified. The difficulty is commonly compounded by interactions with irrelevant biological features with variable representation across the population, and instrumental variations of technical origin [2, 3]. The challenge, in short, is one human experts can only imperfectly meet, inevitably exhibiting marked variation with skill and experience [4, 5] and therefore training of ML-based models to automate morphological characterisation is intrinsically difficult.
@@
A desirable automated cell characterisation model would have the following five key properties. Firstly, it should be robust to domain shift and generalise to different biological, pathological, and instrumental contexts and class distributions [6, 7]. Secondly, the model should achieve high data efficiency, performing well despite sparse ground truth labels and restricted access to comprehensive datasets as commonly found in clinical applications. Thirdly, the decision of the model should aim to be interpretable where possible, as the reasoning behind a model’s decisions may be as important as the decisions themselves [8, 9]. Fourth, the model should have the ability to identify rare or previously unseen patterns of features as such cases fall outside the model’s competence and must be so highlighted. This is particularly important in clinical applications yet often overlooked in model development and evaluation [10]. Finally, the model should be able to quantify the uncertainty attached to its decision, a feature often neglected in model assessments [11].
Optimally performing discriminative ML classification models are able to approximate human performance at classifying cells into a priori defined morphological classes. However, they are not inherently designed with desirable properties such as tolerating domain shifts, class imbalances, anomalies and complex interactions between the outcomes and biological, pathological, and instrumental features. Numerous studies have reported high accuracy in various medical image analysis tasks [12, 13], but we show that none exhibit the desirable properties we described and these open challenges remain to be addressed, limiting the clinical applicability of such models. Therefore, we introduce CytoDiffusion, a modelling approach whose primary aim is not simply to approximate a human expert in a more cost-effective, reproducible, and scalable way, but to capture the space of possible morphological appearances with high-fidelity using a diffusion-based generative model paired with a novel classification approach.
CytoDiffusion is developed and applied to real-world clinical challenges, building on recent work [14, 15, 16, 17] using generative models for classification challenges. A key advantage of generative classifiers is their ability to model the full data distribution rather than merely a decision boundary [14] and these have shown promising robustness to distribution shifts and ability to avoid shortcut solutions. In the context of blood cell image classification, CytoDiffusion is compelled to learn the complete morphological characteristics of each cell type, rather than focusing on easily identified but potentially unreliable features. Figure 1 illustrates the proposed modelling approach and the contributions of this work are: (a) a novel application of latent diffusion models for blood cell image classification, (b) an evaluation framework that goes beyond accuracy and other standard metrics, incorporating domain shift robustness, anomaly detection capability, and performance in low-data regimes, (c) a new dataset of blood cell images that includes artefacts and labeller confidence scores, addressing key limitations in existing datasets, (d) a principled framework for the evaluation of model and human confidence based on established psychometric modelling techniques and (e) a method for generating interpretable heatmaps to explain model’s decisions.
@@
Through these contributions, we aim to establish a new standard for the development and assessment of blood cell image classification models. Our work addresses several important aspects of clinical applicability, including robustness, interpretability, and reliability. We propose that the research community adopt these evaluation tasks and metrics when assessing new models for blood cell image classification. By going beyond simple discriminative statistics, and other conventional benchmarks, we can develop models that are not only high-performing but also trustworthy and clinically relevant.
Results
@@
We begin by validating the quality of images generated by CytoDiffusion through an authenticity test. Next, we assess the model’s in-domain performance on standard classification tasks across multiple datasets. We then examine CytoDiffusion’s ability to quantify uncertainty, comparing its metacognitive capabilities with those of human experts. Following this, we evaluate the model’s proficiency in anomaly detection, crucial for identifying rare or unseen cell types. We proceed to test the model’s robustness to domain shifts, simulating real-world variability in imaging conditions. Subsequently, we investigate CytoDiffusion’s efficiency in low-data regimes, a critical consideration for medical applications where large, well-annotated datasets may be scarce. Finally, we demonstrate CytoDiffusion’s explainability through the generation of counterfactual heatmaps, providing interpretable insights into its decision-making process.
CytoDiffusion Generates Images Indistinguishable From Real Images
@@
To assess whether our model could effectively capture the distribution of blood cell images, we designed an authenticity test. CytoDiffusion was trained on a dataset comprising 32,619 images. The results of the test, which involved expert haematologists assessing the authenticity of the images, underscore the model’s capability to replicate the complex morphological characteristics of blood cells. The ten haematologists participating in the study evaluated 288 images each, resulting in 2,880 individual assessments. They demonstrated an overall accuracy of 52.3%, 95% CI: [50.5%, 54.2%] in distinguishing between real and synthetic images, with a sensitivity of 55.8% and a specificity of 48.9%. This performance is comparable to random guessing, indicating that the synthetic images produced by CytoDiffusion are almost indistinguishable from real blood cell images, even to experienced professionals.
@@
In addition to the authenticity test, we evaluated the quality of conditional synthesis for each blood cell type. The haematologists’ classification of the synthetic images aligned with the model’s intended cell type with an agreement rate of 98.6%. Overall, these results affirm that CytoDiffusion not only generates high-quality synthetic images but also accurately captures the class-defining morphological features of blood cells (see Supplementary Figure 8).
CytoDiffusion Demonstrates Competitive Classification Performance
@@
To establish a baseline for our diffusion-based classifier, we evaluated CytoDiffusion’s performance on standard in-domain classification tasks. We used three datasets for this evaluation: CytoData (our custom dataset), Raabin-WBC [18], and PBC [19] (see Methods for details). Table 1 presents a comparison of the performance of our model against other competitor methods for these datasets. CytoDiffusion demonstrates competitive performance across all datasets, particularly in balanced accuracy. This metric is especially relevant given the significant class imbalances present in these datasets. For instance, CytoData’s class distribution ranges from 103 to 1425 images per class, while Raabin-WBC’s test set contains between 89 to 2660 images per class.
@@
Raabin-WBC Test-A Tavakoli et al. [20] 2021 94.65 92.76
@@
Table 1: Comparison of competitor methods on the CytoData, Raabin-WBC (Test-A) and the PBC dataset. Best results are in bold, second-best are underlined. Balanced accuracy is reported where available.
CytoDiffusion Exhibits Superhuman Uncertainty Quantification
@@
Our dataset uniquely incorporates human expert confidence for all images, providing a rare opportunity to compare model uncertainty with human expert uncertainty. Quantifying uncertainty is complicated by two cardinal aspects of the task. First, uncertainty has no reliable ground truth in real-world settings. Second, uncertainty in the present context contains an aleatoric component–the constitutional discriminability of the classes–and an epistemic component–the agent’s ability to discriminate between them. The former is determined by the domain, the latter by the characteristics of the agent. In the ideal case, the epistemic uncertainty is zero, and the agent’s uncertainty is wholly aleatoric, determined by how much discriminant signal the data contains. In such a case, the relation between uncertainty and accuracy ought to approximate that of an ideal psychophysical observer detecting a noisy signal, i.e. the uncertainty measure should resemble discriminability.
This insight allows us to deploy the mature conceptual apparatus of psychometric function estimation to the task of evaluating an agent’s uncertainty. In brief, an ideal agent should exhibit a sigmoidal relationship between confidence–the inverse of uncertainty–and accuracy, compactly described by four parameters. The threshold of the function is its location at a chosen accuracy level; the slope is the width or slope of the function; the floor is the chance level, determined by the task; and the ceiling is the lapse level, determined by the error rate where discriminability is maximally easy. Here we employ well-established Bayesian psychometric modelling techniques to derive a psychometric function for CytoDiffusion’s performance (Figure  2A), revealing an excellent fit, with tight posterior distributions on the key threshold and width parameters (Figure  2A, inset axes). Although direct measurement is impossible, this suggests CytoDiffusion’s uncertainty is dominated by the aleatoric component and its behaviour is close to that of an ideal observer.
This conclusion is reinforced by evaluating individual human expert performance, judged against expert consensus, with CytoDiffusion’s confidence as the measure of discriminability. The resultant function, illustrated for Expert 2 (Figure  2B), not only exhibits a good fit, but describes the relationship better than consensus human expert confidence (Figure  2C), suggesting CytoDiffusion’s metacognitive abilities are superior to human experts here. Examination of the estimated threshold and width parameters for each human expert, with CytoDiffusion (Figure  2D)) or human expert (Figure  2E) confidence, shows CytoDiffusion’s measure can distinguish between the varying abilities of human experts better than they themselves can.
@@
Figure 2: Bayesian psychometric analysis of model and expert performance. A. Psychometric function of CytoDiffusion’s performance on our custom dataset with model confidence as the index of discriminability. The 95 per cent credibility interval of the threshold of the function, estimated at the point of 80 per cent correct unscaled by lapse and guess rate, is shown as an error bar, and the posterior distributions of threshold and width are presented in the inset axes. Note tight bounds on the parameters. B. Psychometric function of the performance of a single human expert, derived from a randomly selected subset of 200 cases for which a consensus ground truth is available, with model confidence as the index of discriminability. C. Psychometric function of the expert shown in B, with consensus expert confidence as the index of discriminability. D & E. Width and threshold parameter estimates and their 95 per cent credibility intervals for each of the six human experts with model confidence (D) and consensus expert confidence (E) as the indices of signal strength.
CytoDiffusion Excels at Detecting Anomalous Cell Types
@@
For each dataset, we state the anomalous class and the performance, given by the area under the received operating characteristic curve (AUC), of CytoDiffusion in detecting it. For the Bodzas dataset [35], with blasts as the abnormal class, ViT achieved an AUC of 0.698 whereas CytoDiffusion achieved an AUC of 0.982. Similarly, for the PBC dataset [19], with erythroblasts as abnormal, ViT achieved an AUC of 0.918 whereas CytoDiffusion achieved an AUC of 0.990. These high AUC values demonstrate our model’s capability to distinguish between normal cells it was trained on and abnormal cell types not present in the training data.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionCertainty00.20.40.60.8100.20.40.60.81PBCBodzas
@@
Figure 3: Kernel density estimate figures comparing the anomaly detection performance of ViT-B/16 (top row) with CytoDiffusion (bottom row) for blasts (left column) and erythroblasts (right column). The horizontal axis represents the certainty measure, normalised to $[0,1]$ by dividing by the maximum certainty for each model and dataset. The AUC values indicate the model’s ability to distinguish between normal and abnormal samples.
CytoDiffusion Shows Robustness to Domain Shifts
To assess the generalisability of the CytoDiffusion model, it was trained using the Raabin-WBC dataset, with testing performed on two distinct datasets: Test-B and LISC. Test-B is part of the Raabin-WBC dataset, but the authors have provided a specific split for testing purposes. Both Test-B and LISC differ from the training data in camera and microscope types, while LISC additionally uses a different staining technique. In Table 2, we see that CytoDiffusion outperforms the competitor models particularly in LISC where there is a gain of 11.47 in balanced accuracy against the next best model with no overlap of confidence intervals.
Method Test-B LISC
@@
Table 2: Performance under domain shift conditions. Models were trained on the Raabin-WBC dataset and evaluated on Test-B and LISC datasets, representing different degrees of domain shift. Balanced accuracy scores (mean $\pm$ standard deviation) from five independent training sessions are reported.
CytoDiffusion Outperforms Competitors in Low-Data Scenarios
To compare model performances under different training data availability rates, we used the PBC dataset [19]. with training carried out with subsets of 10, 20, 50, and 150 images per class, simulating conditions of sparse data availability. In Figure 4 we see that CytoDiffusion consistently outperforms competitor discriminative models EfficientNetV2-M and ViT-B/16 across all data availability levels, with the advantage particularly pronounced in the most data-scarce conditions.
@@
Figure 4: Model performance comparison under low-data conditions. Shaded areas represent the standard deviation from five independent training sessions.
CytoDiffusion Provides Visual Explanations Through Counterfactual Heatmaps
@@
Counterfactual heatmaps highlight the regions of an image that would need to change for it to be classified as a different cell type. In Figure 5, we used an eosinophil as an example and prompted the model to consider what alterations would be necessary for this cell to be classified as a neutrophil, generating a heatmap ($\boldsymbol{H}_{\text{neutrophil}}$) that highlights regions where there is large error in the latent space between the two classes. The overlay of this heatmap on the original image reveals that the model focuses primarily on distinguishing granularity between neutrophils and eosinophils, with areas of significant colour deviation from the background indicating the most critical regions of difference.
Original Heatmap Overlay
@@
Refer to caption “If this image were a neutrophil, what would you change?” Refer to caption Refer to caption
@@
Figure 5: Counterfactual heatmap visualisation. Left: Original eosinophil image. Centre-right: Counterfactual heatmap ($\boldsymbol{H}_{\text{neutrophil}}$) illustrating the areas that differ when the model considers the image as a neutrophil. (For an example of how a neutrophil can look, see Figure 6.) Far right: Overlay of the thresholded heatmap on the original image.
@@
To provide a comprehensive view of CytoDiffusion’s abilities across all cell types, in Figure 6) we generated counterfactual heatmaps for each possible class transition in the PBC dataset. This visualisation offers insights into the model’s decision-making process for each cell type. For instance, when considering the transition from neutrophil to eosinophil (row 2, column 7), the model highlights regions in the cytoplasm (darker areas) where features should be added, while largely maintaining the nuclear shape.